# Unveiling the optimization process of Physics Informed Neural Networks: How accurate and competitive can PINNs be?

### v1

@@ 1 Introduction
Recent advancements in Physics-Informed Neural Networks (PINNs) have positioned them as serious contenders in the domain of computational physics, disrupting the longstanding monopoly held by classical numerical methods. This disruptive potential arises from their innate ability to integrate domain-specific physics principles with the powerful learning capabilities of neural networks. However, despite their large potential, the nascent nature of this field has left significant gaps in the literature surrounding PINNs. The lack of rigorous mathematical analysis and specificity in methodologies often obscures crucial insights into their limitations, both in terms of accuracy and efficiency.
@@
While PINNs exhibit remarkable adaptability to complex physical systems and have demonstrated interesting results across various domains, their efficacy relies on several factors that warrant closer scrutiny. One critical aspect is the architectural design of the neural network, which influences its ability to capture intricate physical phenomena accurately. Additionally, the choice of loss functions and optimization techniques also play a pivotal role in the overall performance of PINNs. Furthermore, the computational overhead associated with training large-scale PINNs raises questions about their scalability and efficiency, particularly in comparison to established numerical methods that have been optimized over decades.
@@
Addressing these challenges needs a concerted effort from researchers to establish comprehensive benchmarks, standardized evaluation metrics, and theoretical frameworks that elucidate the fundamental principles governing the behavior of PINNs. Only through rigorous analysis and systematic experimentation can we unlock the full potential of PINNs and turn them into reliable tools for solving complex physical problems across diverse fields, ranging from fluid dynamics and solid mechanics to quantum physics and beyond.
@@
Having recognized the need to move beyond the trial-and-error approaches prevalent in most of the PINNs literature, this paper aims to improve our understanding of the fundamental aspects that define their performance. At its core, PINNs training reduces to an optimization problem, prompting us to revisit the basics of optimization theory. We explore some intricacies of PINN optimization, seeking to identify the bottleneck that hinders their precision in various physical applications. Our focus extends beyond mere adjustments in network size, architecture, activation functions, or other hyperparameters. Instead, we argue that perhaps the most important ingredients are the fundamental principles that govern the optimization process. Through the exploration of different techniques, we bracket the boundaries of precision achievable with PINNs across diverse physical scenarios, focusing on the pivotal role played by the choice of the optimizer in determining the accuracy and efficiency of PINN solutions.
@@
The paper is structured as follows: Section 2 provides a brief overview of our PINN framework and discusses key issues related to commonly used optimizers. In Section 3, we thoroughly examine a relatively simple case to demonstrate the significant impact of selecting the appropriate optimizer and how such a choice can minimize network size while achieving excellent results, surpassing previous studies with similar problems. Additionally, in Section 4, we present a comprehensive set of physical problems spanning various fields, illustrating that the insights from the preceding section can be extended to a large variety of problems. Finally, Section 5 summarizes our findings and outlines potential paths for future improvements.
@@
$\mathcal{J}=\frac{1}{N}\sum_{i=1}^{N}||\mathcal{L}u(x_{\alpha i};\bm{\Theta})-G(x_{\alpha i},u(x_{\alpha i};\bm{\Theta}))||^{2}~{}.$ (2)
To completely describe the physical system, we also need to impose boundary conditions on a boundary $\partial D$. When Dirichlet and periodic boundary conditions are involved, we impose them through the so-called hard-enforcement [1, 3, 4, 5]. This means that we redefine the solution so that they are satisfied by construction independently of the PINN’s output. In the case of Dirichlet boundary conditions this can be achieved by redefining the solution in the following way:
@@
where $f_{b}$ is any smooth function that satisfies the Dirichlet boundary conditions when $x_{\alpha}\in\partial D$, $h_{b}$ is a smooth function that is equal to zero when $x_{\alpha}\in\partial D$, and $\mathcal{N}$ is the output of the PINN.
Periodic boundary conditions can also be readily hard-enforced. For example, let us designate $x_{\beta}$ and $x_{\gamma}$ as subsets of $\partial D$ where we apply Dirichlet and periodic boundary conditions, respectively. Then, we redefine the solution as follows:
$u(x_{\alpha};\bm{\Theta})=f_{b}(x_{\beta})+h_{b}(x_{\beta})\mathcal{N}(x_{\beta},\cos{(x_{\gamma})},\sin{(x_{\gamma})};\bm{\Theta}).$ (4)
One can indeed check that, $u$ has the desired behaviour at all the borders: Dirichlet boundary conditions are satisfied because $u$ is equal to $f_{b}$ at the borders of $x_{\beta}$ and periodic boundary conditions are satisfied because the coordinates $x_{\gamma}$ are first passed to the Fourier basis.
For Neumann or Robin boundary conditions, hard-enforcement is also possible but not straightforward to implement and can result in rather cumbersome and complicated expressions. In these cases, one can always use the soft-enforcement, i.e. add a term in the loss function (2) that takes into account the residuals of the boundary condition on a sample of points at the boundary.
The loss function $\mathcal{J}$ is a multidimensional scalar function of the parameters $\bm{\Theta}$. Its minimization requires a robust optimization algorithm that updates the parameters after each training epoch. Two very popular choices in the literature of PINNs are the Adam [6] and BFGS [7, 8, 9, 10] optimizers.
The Adam optimizer has consistently been a fundamental component in the training of various machine learning applications and PINNs. However, it has recently become clear that quasi-Newton methods such as BFGS or its low-memory variant L-BFGS [11] can achieve more accurate results in significantly fewer iterations than Adam, but they are more prone to be trapped at saddle points. The state-of-the-art training schemes involve a combination of these two optimizers, using Adam for the initial iterations to handle better the possible presence of saddle points, and then using BFGS / L-BFGS to accelerate convergence.
2.1 Brief review of optimization procedures.
@@
The simplest case is to consider the gradient decent algorithm, where $H_{k}=I$ and $\alpha_{k}$ is equal to a small positive constant, which has linear convergence. Adam’s method can be recovered by using $H_{k}=I$ and a formula for calculating a specific $\alpha_{k}$ for each parameter. It can be seen as a more sophisticated variant of the gradient descent algorithm, which has better convergence properties but is still linear. Newton’s method can be recovered by considering $H_{k}$ to be the exact inverse of the Hessian matrix of $\mathcal{J}$. It is a second-order method that can converge in very few iterations but with a large increase in the computational cost of each iteration: it requires the explicit calculation of second derivatives of the loss function and the inversion of a high dimensional matrix. The so-called Quasi-Newton methods lie in between. They use some approximation of the inverse Hessian that requires only the first derivatives of the loss function and involves only matrix-vector multiplications, resulting in superlinear convergence (not yet quadratic), but are much faster than Newton’s method per iteration. The step size $\alpha_{k}$ is usually chosen with inexact line search procedures that preserve the positive-definiteness of $H_{k}$ by imposing certain restrictions on $\alpha_{k}$ (for example, the Wolfe conditions [13]).
@@
where $\otimes$ denotes the tensor product of two vectors and $\tau_{k},\phi_{k}$ are respectively the scaling and the updating parameters, which in general change between iterations. For $\tau_{k}=1$ and $\phi_{k}=1$, we recover the standard BFGS algorithm. Other choices for the value of these parameters define different methods, as we will explore in detail in the next section.
@@ 3 Case study: neutron star magnetospheres
We now focus on the problem of force-free neutron star magnetospheres in the non-rotating axisymmetric regime as a baseline case. This problem was examined in detail in [16], illustrating the potential of the PINN approach in this particular astrophysical scenario. Here, we revisit this study to highlight the significant influence that the selection of the optimization algorithm can have on performance. While a detailed exposition of the theoretical background can be found in the aforementioned paper, here we provide a brief overview of the core concepts and equations for completeness.
It is convenient to use compactified spherical coordinates $x_{\alpha}=(q,\mu,\phi)$, where $q=1/r$ and $\mu=\cos{\theta}$ and introduce dimensionless units $R$ (radius of the star) and $B_{0}$ (surface magnetic field at the equator of the dipolar component). In axisymmetry, the magnetic field $\bm{B}$ can be described in terms of two poloidal and toroidal stream functions $\mathcal{P}$ and $\mathcal{T}$ as
where $\bm{e}_{\phi}$ is the unit vector in the $\phi$ direction.
$\nabla\mathcal{P}\times\nabla\mathcal{T}=0,$ (12)
which implies that $\mathcal{T}=\mathcal{T}(\mathcal{P})$, and the so-called Grad-Shafranov equation
$\triangle_{\mathrm{GS}}\mathcal{P}+\mathcal{T}\frac{d\mathcal{T}}{d\mathcal{P}}=0~{}.$ (13)
@@ 3.1.1 Impact of the optimization algorithm
The architecture and training hyperparameters that we use are outlined in table 1. We address the problem utilizing four different optimization updates: Adam, BFGS, self-scaled BFGS (SSBFGS) and self-scaled Broyden (SSBroyden). The first two are usually employed in most PINN applications. The last two methods are variations of equation (10) tailored with specific selections of the parameters $\tau_{k}$ and $\phi_{k}$. These methods are well-established in optimization theory and have demonstrated advantages. In fact, they could be considered as modifications to the BFGS formula rather than new, standalone optimizers. The term ”self-scaled” means that $\tau_{k}\neq 1$ and corresponds to the usual BFGS formula, but with a scaling factor multiplying the approximation $H_{k}$ of the inverse Hessian, while “Broyden” method assumes $\phi_{k}\neq 1$. For SSBFGS we use the choices suggested in [17]:
$\phi_{k}$ $=1.$ (19)
In A we elaborate on the details of this optimizer and show how $\tau_{k}^{(1)}$ can be efficiently calculated without the need to invert the matrix $H_{k}$. For SSBroyden we use the choices suggested in [15]:
$\tau_{k}^{(2)}$ =$=\begin{cases}\tau_{k}^{(1)}\min\left(\sigma_{k}^{-1/(n-1)},\frac{1}{\theta_{k}^{(1)}}\right)\mathrm{if}\theta_{k}>0\\ \min\left(\tau_{k}^{(1)}\sigma_{k}^{-1/(n-1)},\sigma_{k}\right)\mathrm{if}\theta_{k}\leq 0,\end{cases}$ (20)
$\phi_{k}^{(1)}$ $=\frac{1-\theta_{k}}{1+a_{k}\theta_{k}},$ (21)
where $n=\mathrm{size}({\bm{\Theta}}_{k})$ is the total number of the trainable parameters and $\sigma_{k},\theta_{k},a_{k}$ are intermediate auxiliary variables, whose definitions are given in B. We should stress here that $\tau_{k}$ and $\phi_{k}$ should respect certain restrictions in order to ensure global and super-linear convergence of the updating algorithm. Furthermore, the choices presented here are not the only possible ones. We extensively experimented with different options and determined that these choices consistently led to improved convergence and more precise solutions across the diverse range of problems we explored.
Figure 1 shows the evolution of the loss function with the number of iterations for the four optimizers considered. In all cases, we use Adam for the initial training phase in order to avoid possible saddle points and get closer to a global minimum before accelerating convergence with a quasi-Newton method. The impact on the convergence of the quasi-Newton formulae is glaring when compared to Adam. BFGS achieves a loss function that is six orders of magnitude smaller than Adam, which is reduced by a further two and three orders of magnitude by its modifications SSBFGS and SSBroyden respectively. This is further reflected in the absolute and relative errors of $\mathcal{P}$ and the magnetic field components (which depend on the first derivatives of $\mathcal{P}$), as can be seen in table 4. For further insight into the error estimates, the interested reader can consult all pertinent details in C.
The established interpretation for this phenomenon attributes it to a poorly scaled loss function $\mathcal{J}$, which results in an ill-conditioned Hessian matrix $\mathrm{hess}(\mathcal{J})$ close to the minimum [18, 19, 20]. This characteristic is intrinsic to PINN configurations and is associated with formulating the loss function based on a differential operator [21, 22].
An approach to grasp this phenomenon is through an examination of the spectrum of the Hessian matrix. As depicted in Figure 2, $\mathrm{hess}(\mathcal{J})$ demonstrates a broad eigenvalue spectrum, with numerous eigenvalues closely approaching zero and some outliers displaying larger magnitudes. Additionally, the condition number, represented by
@@
Another way of understanding the influence of quasi-Newton methods, as pointed out in [21, 22], is through the preconditioning of the loss function. At every iteration step, the inverse Hessian approximation acts as a preconditioner that maps the parameter space $\bm{\Theta}$, where the Hessian is ill-conditioned, to a new space $\bm{z}=H_{k}^{-1/2}\bm{\Theta}$, where the conditioning is much better. The rate of convergence of the different quasi-Newton methods will be strongly affected by the conditioning of the Hessian in this new space. To observe this effect mathematically, consider that representing equation (5) in the $\bm{z}$ space, corresponds to the usual formula of a gradient-descent method
@@
Please note that while the exact Newton’s formula theoretically promises perfect conditioning, ensuring uniform gradients in all directions, its practical implementation poses numerical challenges. This is due to the necessity of solving an ill-conditioned linear system of equations to determine the Newton descent direction. Typically, such ill-conditioning arises from slight variations in matrix coefficients, particularly in the Hessian, resulting in significant deviations in the solution accuracy ([25]).
To appreciate the effect of the Hessian conditioning, figure 3 shows the spectra of the Hessian expressed in the $\bm{z}$ space for each optimization algorithm. Each of these histograms is computed by letting the PINN to be trained additional iterations with a fixed training set until we arrive to similar values of $\mathcal{J}$, which we set to be very low ($\mathcal{J}\sim 10^{-13}$) in order to be close to the minimum. As we can observe, the standard BFGS algorithm gives an ill-conditioned Hessian in the $\bm{z}$ space close to the minimum, whereas with the BFGS modifications we obtain better-conditioned matrices. This is reflected in the number of iterations needed for each algorithm to arrive at the value of loss prescribed above: the BFGS algorithm needed $\sim 40000$ iterations, whereas the BFGS modifications only needed $\sim 1000$.
Figure 1: Loss function vs. iterations for the current-free Grad-Shafranov equation. Results are obtained for the four optimization algorithms considered in section 3.1.1.
Figure 2: Eigenvalue spectrum of $\mathrm{hess}\left(\mathcal{J}\left(\bm{\Theta}\right)\right)$. It is obtained after the loss function has been reduced to $\mathcal{J}\sim 10^{-13}$ for a fixed set of points in order to be as close to the minimum as possible.
Figure 3: Eigenvalue spectra of the preconditioned Hessian matrix $\mathrm{hess}\left(\mathcal{J}\bm{z}\right)$ for BFGS (3(a)), SSBFGS (3(b)) and SSBroyden (3(c)).

### v2

@@ 1 Introduction
Recent advancements in Physics-Informed Neural Networks (PINNs) have positioned them as serious contenders in the domain of computational physics [1, 2], disrupting the longstanding monopoly held by classical numerical methods in many and varied physical applications, such as fluid and solid mechanics [3, 4], quantum optics [5], black-hole spectroscopy [6], or radiative transfer [7], among many others. This disruptive potential arises from their innate ability to integrate domain-specific physics principles with the powerful learning capabilities of neural networks.
However, despite their significant potential, the nascent literature on PINNs is less than a decade old -which is minimal compared to numerical analysis, for example- and more often than desirable it suffers from a relative lack of rigorous mathematical analysis and specificity. The frequent reliance on trial-and-error methodologies, acquired from the machine learning literature, tends to obscure important insights into the origin of mathematical limitations, particularly concerning accuracy and efficiency. Fortunately, as the field matures, an increasing number of studies are delving deeper into the mathematics underlying the components of PINNs.
One of the most challenging aspects of neural networks is the inherently non-convex nature of their optimization problems. As a result, an increasing number of studies are delving deeper into their learning dynamics (see for example [8, 9, 10, 11]). Focusing specifically on PINNs, [12] analyses the convergence process using gradient-descent methods by examining the eigenvalue spectra of the Neural Tangent Kernel (NTK). The NTK has been discussed in many subsequent studies to provide a better understanding of the dynamics of the training process in many physical problems under different modifications of the standard PINN algorithm [13, 14, 15, 16]. The learning dynamics of PINNs has recently been studied in [17] through the lens of the gradient signal-to-noise ratio (GSNR), which let the authors characterize the optimization process of PINNs in different phases, as predicted by the Information Bottleneck theory [18]. As a complementary path, other works in the PINN literature have focused more on the development of new optimization techniques. In [19] they analyzed the behavior of gradient-descent based optimization methods, concluding that training with these algorithms is limited because of the ill-conditioning nature of the parameter space, and suggested various strategies to precondition this space. [20] introduced a novel second-order optimizer that significantly improves the solution returned by L-BFGS for some problems. Learnable optimization, also known in the field of deep learning as learning to learn [21] has also been applied recently to PINNs in [22], improving the results obtained with standard machine learning optimizers.
@@
While PINNs exhibit remarkable adaptability to complex physical systems and have demonstrated interesting results across various domains, their efficacy relies on several factors that warrant closer scrutiny. One critical aspect is the architectural design of the neural network [23, 24, 25, 26], which influences its ability to capture intricate physical phenomena accurately. Other ideas proposed to enhance the performance of PINNs are for example domain decompositions [27], trainable activation functions [28], loss function redefinitions [29], or curriculum strategies [30], among others.
Despite their flexibility, training neural networks can be computationally expensive for many problems. The challenge also makes it difficult to determine which hyperparameters are crucial for improving results. Recent research has focused on this issue, particularly on the weights controlling the residual and supervised terms, with new methods proposed to mitigate the negative effects of their magnitude disparity [31, 32, 33]. Some studies have also tried to give estimates of the errors as a function of different hyperparameters, such as the network size or the number of training points, for particular PDE problems [34, 35, 36, 37, 38, 39]. Scaling PINNs to large-scale or high-dimensional problems can be challenging. The neural network may require more layers or neurons, increasing the risk of overfitting or making the training process unstable. Tackling the so-called Curse of Dimensionality is not a trivial task, and it has been recently studied in [40].
@@
Addressing these and other challenges needs a concerted effort from researchers to establish comprehensive benchmarks, standardized evaluation metrics, and theoretical frameworks that elucidate the fundamental principles governing the behavior of PINNs. Only through rigorous analysis and systematic experimentation can we unlock their full potential and turn them into reliable tools for solving complex physical problems across diverse fields, ranging from fluid dynamics and solid mechanics to quantum physics and beyond.
Although PINNs also have great potential in solving inverse problems (with applications in many and varied scenarios, such as fluid mechanics [41, 42, 43, 44], materials science [45, 46], plasma physics [47], among many others [48, 49, 50]), which are crucial for many scientific and engineering applications, in this work, we focus on the use of PINNs for forward problems, specifically in solving partial differential equations for physics-related applications. Nevertheless, we think that our findings could be also useful in other contexts, including inverse problems, which deserve a more thorough exploration.
@@
This paper aims to improve our understanding of the fundamental aspects that define PINNs performance. At its core, training reduces to an optimization problem, prompting us to revisit the basics of optimization theory. We explore some intricacies of PINN optimization, seeking to identify the bottleneck that hinders their precision in various physical applications. Our focus extends beyond mere adjustments in network size, architecture, activation functions, or other hyperparameters. Instead, we argue that perhaps the most important ingredients are the fundamental principles that govern the optimization process. Through the exploration of different techniques, we bracket the boundaries of precision achievable with PINNs across diverse physical scenarios, focusing on the pivotal role played by the choice of the optimizer in determining the accuracy and efficiency of PINN solutions.
@@
The paper is structured as follows: Section 2 provides a brief overview of our PINN framework and discusses key issues related to commonly used optimizers. We dive into the details behind the optimization process in PINNs and we present our proposed contributions to improve it in Section 3. In Section 4, we thoroughly examine a relatively simple case to demonstrate the significant impact of selecting the appropriate optimizer and how such a choice can minimize network size while achieving excellent results, surpassing previous studies with similar problems. Additionally, in Section 5, we present a comprehensive set of physical problems spanning various fields, illustrating that the insights from the preceding section can be extended to a large variety of problems. Finally, Section 6 summarizes our findings and outlines potential paths for future improvements.
@@
$\mathcal{J}=\frac{1}{N}\sum_{i=1}^{N}|\mathcal{L}u(x_{\alpha i};\bm{\Theta})-G(x_{\alpha i},u(x_{\alpha i};\bm{\Theta}))|^{2}~{}.$ (2)
To completely describe the physical system, we must impose boundary conditions on a boundary $\partial D$. When Dirichlet and periodic boundary conditions are involved, we impose them through the so-called hard-enforcement [51, 52, 53, 54, 55]. This means that we redefine the solution so that they are satisfied by construction independently of the PINN’s output. In the case of Dirichlet boundary conditions, this can be achieved by redefining the solution in the following way:
@@
where $f_{b}$ is a suitable smooth function that satisfies the Dirichlet boundary conditions when $x_{\alpha}\in\partial D$, $h_{b}$ is a suitable smooth function that vanishes when $x_{\alpha}\in\partial D$, and $\mathcal{N}$ is the output of the PINN.
Periodic boundary conditions with periodicity $L$ can also be hard-enforced, as amply discussed in [56]. For example, let us designate $x_{\beta}$ and $x_{\gamma}$ as subsets of $\partial D$ where we apply Dirichlet and periodic boundary conditions, respectively. Then, if we redefine the solution as follows:
$u(x_{\alpha};\bm{\Theta})$ $=f_{b}(x_{\beta})+h_{b}(x_{\beta})~{}\mathcal{N}(x_{\beta},\cos\left({kx_{\gamma}}\right),\sin\left({kx_{\gamma}}\right);\bm{\Theta}),$ (4)
where $k=2\pi/L$, one can indeed check that $u$ has the desired behaviour at all the borders. This parametrization enforces periodicity for $u$ and all its derivatives with respect to $x_{\gamma}$, provided that $f_{b}$ and all its derivatives are also periodic. However, this does not necessarily work for all cases (for example, if we want a periodic function but not its derivatives). For some problems, the full Fourier expansion is formally needed to compute the solution (see e.g. [57, 24, 58]). Another possible approach is to consider a different set of functions, such as Hermite-based interpolation polynomials [56], chosen to be periodic up to the desired order. These polynomials contain additional trainable parameters to enforce the appropriate relation between the function and the derivatives at the boundaries.
For Neumann or Robin boundary conditions, hard-enforcement is also possible but not straightforward to implement and can result in rather cumbersome and complicated expressions. In these cases, one can always use the soft-enforcement, i.e. add terms in the loss function (2) that take into account the residuals of the boundary condition on a sample of points at the boundary. The soft-enforcement approach is highly flexible and can be applied to all the previously discussed cases. However, its main drawback is that its performance depends on selecting additional hyperparameters, such as the number of boundary points and the relative weights assigned to these terms, which influence the focus on boundary conditions during the optimization process.
The loss function $\mathcal{J}$ is a multidimensional scalar function of the parameters $\bm{\Theta}$. Its minimization requires a robust optimization algorithm that updates the parameters after each training iteration. Two very popular choices in the literature of PINNs are the Adam [59] and BFGS [60, 61, 62, 63] optimizers. The Adam optimizer has consistently been a fundamental component in the training of various machine learning applications and PINNs. However, it has recently become clear that quasi-Newton methods such as BFGS or its low-memory variant L-BFGS [64] can achieve more accurate results in significantly fewer iterations than Adam, but they are more prone to be trapped at saddle points. The state-of-the-art training schemes involve a combination of these two optimizers, using Adam for the initial iterations to handle better the possible presence of saddle points, and then using BFGS / L-BFGS to accelerate convergence.
3 Optimization method
3.1 Brief review of optimization procedures
@@
The simplest case is to consider the gradient decent algorithm, where $H_{k}=I$ and $\alpha_{k}$ is equal to a small positive constant, which has linear convergence. The Adam optimizer can be recovered by using $H_{k}=I$ and a formula for calculating a specific $\alpha_{k}$ for each parameter. It can be seen as a more sophisticated variant of the gradient descent algorithm, which has better convergence properties but is still linear. Newton’s method can be recovered by considering $H_{k}$ to be the exact inverse of the Hessian matrix of $\mathcal{J}$. It is a second-order method that can converge in very few iterations but with a large increase in the computational cost of each iteration: it requires the explicit calculation of second derivatives of the loss function to calculate the hessian matrix, and then also its inversion, which can be prohibitively costly in large-scale optimization problems. Moreover, in non-convex problems the Newton’s method could give away from the minimum non-descent directions, ultimately leading to line search failure. The so-called Quasi-Newton methods lie in between. They use some approximation of the inverse Hessian that requires only the first derivatives of the loss function and involves only matrix-vector multiplications, resulting in superlinear convergence (not yet quadratic), but are much faster than Newton’s method per iteration. The step size $\alpha_{k}$ is usually chosen with inexact line search procedures that preserve the positive-definiteness of $H_{k}$ by imposing certain restrictions on $\alpha_{k}$ (for example, the Wolfe conditions [66]).
@@
where $\otimes$ denotes the tensor product of two vectors and $\tau_{k},\phi_{k}$ are respectively the scaling and the updating parameters, which in general change between iterations. For $\tau_{k}=1$ and $\phi_{k}=1$, we recover the standard BFGS algorithm.
3.2 Modifications of optimization algorithm
In this work, we introduce two methods, which we label as self-scaled BFGS (SSBFGS) and self-scaled Broyden (SSBroyden) method respectively. These methods are well-established in optimization theory and have demonstrated certain advantages [68, 69]. In fact, they could be considered as modifications to the BFGS formula rather than new, standalone optimizers. The term “self-scaled” means that $\tau_{k}\neq 1$ and corresponds to the usual BFGS formula, but with a scaling factor multiplying the approximation $H_{k}$ of the inverse Hessian, while “Broyden” method assumes $\phi_{k}\neq 1$.
For SSBFGS we use the choices suggested in [69]:
$\phi_{k}$ $=1.$ (12)
In B we elaborate on the details of this optimizer and show how $\tau_{k}^{(1)}$ can be efficiently calculated without the need to invert the matrix $H_{k}$.
For SSBroyden we use the choices suggested in [68]:
$\tau_{k}^{(2)}$ =$=\begin{cases}\tau_{k}^{(1)}\min\left(\sigma_{k}^{-1/(n-1)},\frac{1}{\theta_{k}}\right)&\quad\mathrm{if}~{}\theta_{k}>0\\ \min\left(\tau_{k}^{(1)}\sigma_{k}^{-1/(n-1)},\sigma_{k}\right)&\quad\mathrm{if}~{}\theta_{k}\leq 0,\end{cases}$ (13)
$\phi_{k}^{(1)}$ $=\frac{1-\theta_{k}}{1+a_{k}\theta_{k}},$ (14)
where $n=\mathrm{size}({\bm{\Theta}}_{k})$ is the total number of the trainable parameters and $\sigma_{k},\theta_{k},a_{k}$ are intermediate auxiliary variables, defined through the following relations:
$h_{k}$ $=\frac{\bm{y}_{k}\cdot H_{k}\bm{y}_{k}}{\bm{y}_{k}\cdot\bm{s}_{k}},$ (16)
$a_{k}$ $=h_{k}b_{k}-1$ (17)
$c_{k}$ $=\sqrt{\frac{a_{k}}{a_{k}+1}},$ (18)
$\rho_{k}^{-}$ $=\min\left(1,h_{k}\left(1-c_{k}\right)\right),$ (19)
$\theta_{k}^{-}$ $=\frac{\rho_{k}^{-}-1}{a_{k}},$ (20)
$\theta_{k}^{+}$ $=\frac{1}{\rho_{k}^{-}},$ (21)
$\theta_{k}$ $=\max\left(\theta_{k}^{-},\min\left(\theta_{k}^{+},{\frac{1-b_{k}}{b_{k}}}\right)\right)$ (22)
$\sigma_{k}$ $=1+a_{k}\theta_{k}.$ (23)
We should stress here that $\tau_{k}$ should respect certain restrictions in order to ensure global and super-linear convergence of the updating algorithm. In particular, for any $\theta_{k}$ that satisfies the inequality $(1-\nu_{1})(-\frac{1}{a_{k}})\leq\theta_{k}\leq\nu_{2}<\infty$, there exists a $\tau_{k}$ for which the inequalities $(1-\nu_{1})(-\frac{1}{a_{k}})\leq\tau_{k}\theta_{k}\leq 1-\nu_{3},\nu_{4}\leq\tau_{k}\leq 1$ hold, with $\nu_{i}>0$ constants associated with the machine accuracy [70]. Furthermore, the choices presented here are not the only possible ones. We extensively experimented with different options and determined that these choices consistently led to improved convergence and more precise solutions across the diverse range of problems we explored. For a comprehensive review with numerous references regarding diverse quasi-Newton methods within the self-scaled Broyden family, see [71].
3.3 Modifications of the loss function
Apart from improvements in the optimization algorithm, we also investigate the effect of using a slightly modified version of the usual MSE loss function (2). In particular, we explore the consequences of evaluating the loss function through a user-defined monotonically increasing function $g$:
$\mathcal{J}_{g}=g\left(\mathcal{J}\right).$ (24)
Near the minimum, the second term can be neglected, because $\nabla\mathcal{J}\simeq 0$, and both Hessian matrices are proportionally related to each other with ratio $g^{\prime}(\mathcal{J})$. Two obvious choices for $\mathcal{J}_{g}$ are
$\mathcal{J}_{1/2}$ $\equiv\sqrt{\mathcal{J}},$ (26)
$\mathcal{J}_{\log}$ $\equiv\log{\mathcal{J}}.$ (27)
Since both the square root and the natural logarithm exhibit derivatives exceeding 1 when $\mathcal{J}\ll 1$, they have the potential to accelerate convergence.
@@ 4 Case study: neutron star magnetospheres
We now focus on the problem of force-free neutron star magnetospheres in the non-rotating axisymmetric regime as a baseline case. This problem was examined in detail in [72], illustrating the potential of the PINN approach in this particular astrophysical scenario. Here, we revisit this study to highlight the significant influence that the selection of the optimization algorithm and/or loss function can have on performance. While a detailed exposition of the theoretical background can be found in the aforementioned paper, we provide a brief overview of the core concepts and equations in A for completeness.
It is convenient to use compactified spherical coordinates $x_{\alpha}=(q,\mu,\phi)$, where $q=1/r$ and $\mu=\cos{\theta}$ and introduce dimensionless units $R$ (radius of the star) and $B_{0}$ (surface magnetic field at the equator of the dipolar component). In axisymmetry, the magnetic field $\bm{B}$ can be described in terms of two poloidal and toroidal stream functions $\mathcal{P}$ and $\mathcal{T}$. The problem then reduces to the so-called Grad-Shafranov equation
$\triangle_{\mathrm{GS}}\mathcal{P}+\mathcal{T}\frac{d\mathcal{T}}{d\mathcal{P}}=0~{},$ (28)
@@ 4.1.1 Impact of the optimization algorithm
The architecture and training hyperparameters that we use are outlined in table 1. We address the problem utilizing four different optimization methods: Adam, BFGS, SSBFGS and SSBroyden.
We refresh the randomly sampled training points every 500 iterations. We found this simple approach effective, providing a good balance between loss reduction and avoiding overfitting. While more advanced sampling techniques could be used to further improve results (see [73] and references for details), we stick to this simpler method to isolate and focus on the impact of our proposed methodology: without other refinements, variations in optimizers will lead to improvements in precision of several orders of magnitude.
The left panel of figure 1 shows the evolution of the loss function with the number of iterations for the four optimizers considered. In all cases, we use Adam for the initial training phase in order to avoid possible saddle points and get closer to a global minimum before accelerating convergence with a quasi-Newton method. The impact on convergence of the quasi-Newton formulae is glaring when compared to Adam. BFGS achieves a loss function that is six orders of magnitude smaller than Adam, which is reduced by a further two and three orders of magnitude by its modifications SSBFGS and SSBroyden respectively. This is further reflected in the absolute and relative errors of $\mathcal{P}$ and the magnetic field components (which depend on the first derivatives of $\mathcal{P}$), as can be seen in table 2.
The established interpretation of this phenomenon attributes it to a poorly scaled loss function $\mathcal{J}$, which results in an ill-conditioned Hessian matrix $\mathrm{hess}(\mathcal{J})$ close to the minimum [74, 75, 76]. This characteristic is intrinsic to PINN configurations and is associated with formulating the loss function based on a differential operator [77, 20].
One approach to grasp this phenomenon is through the examination of the spectrum of the Hessian matrix. As depicted in the upper left panel of figure 2, $\mathrm{hess}(\mathcal{J})$ demonstrates a broad eigenvalue spectrum, with numerous eigenvalues closely approaching zero and some outliers displaying larger magnitudes. Additionally, the condition number, represented by
@@
Another way of understanding the influence of quasi-Newton methods, as pointed out in [77, 20], is through the preconditioning of the loss function. At every iteration step, the inverse Hessian approximation acts as a preconditioner that maps the parameter space $\bm{\Theta}$, where the Hessian is ill-conditioned, to a new space $\bm{z}=H_{k}^{-1/2}\bm{\Theta}$, where the conditioning is much better. The rate of convergence of the different quasi-Newton methods will be strongly affected by the conditioning of the Hessian in this new space. To observe this effect mathematically, consider that representing equation (5) in the $\bm{z}$ space corresponds to the usual formula of a gradient-descent method
@@
Please note that while the exact Newton’s formula theoretically promises perfect conditioning, ensuring uniform gradients in all directions, its practical implementation poses numerical challenges. This is due to the necessity of solving an ill-conditioned linear system of equations to determine the Newton descent direction. Typically, such ill-conditioning arises from slight variations in matrix coefficients, particularly in the Hessian, resulting in significant deviations in the solution accuracy (see for example [80]).
To appreciate the effect of the Hessian conditioning, the upper right, middle left and lower left panels in figure 2 show the spectra of the Hessian expressed in the $\bm{z}$ space for each optimization algorithm. Each of these histograms is computed by letting the PINN to be trained additional iterations with a fixed training set until we arrive to similar values of $\mathcal{J}$, which we set to be very low ($\mathcal{J}\sim 10^{-13}$) in order to be close to the minimum. As we can observe, the standard BFGS algorithm gives an ill-conditioned Hessian in the $\bm{z}$ space close to the minimum, whereas with the BFGS modifications we obtain better-conditioned matrices. This is reflected in the number of iterations needed for each algorithm to arrive at the value of loss prescribed above: the BFGS algorithm needed $\sim 40000$ iterations, whereas the BFGS modifications only needed $\sim 1000$.
Figure 1: Evolution of the loss function with training iterations for the current-free Grad-Shafranov equation. In the left panel, results are obtained with the four optimization algorithms considered in section 4.1.1. In all cases, the standard MSE loss function $\mathcal{J}$ is used for training and the Adam optimizer is employed for the initial training phase. In the right panel, results are obtained with the loss function modifications considered in section 3.3. In all cases, the Adam optimizer is employed for the initial training phase and the BFGS algorithm for the quasi-Newton stage. Note that the plot shows the MSE loss $\mathcal{J}$ and not $\mathcal{J}_{1/2}$ or $\mathcal{J}_{\mathrm{log}}$ that were used during training.
Figure 2: Eigenvalue spectra of the Hessian matrix $\mathrm{hess}(\mathcal{J}(\bm{z}))$ in the $\bm{z}$-space close to the minimum ($\mathcal{J}\sim 10^{-13}$) for the cases discussed in section 4.1. The upper left panel corresponds to no preconditioning ($H_{k}~{}=~{}I,\bm{z}~{}=~{}\bm{\Theta})$. The upper right panel corresponds to the usual BFGS optimization algorithm and MSE loss function. The middle and lower left panels correspond to the BFGS formula modifications while the middle and lower right panels correspond to the loss function modifications.